# Decentralized Domain Generalization with Style Sharing: Formal Model and Convergence Analysis

### v3

@@ Abstract
Much of federated learning (FL) focuses on settings where local dataset statistics remain the same between training and testing. However, this assumption often does not hold in practice due to distribution shifts, motivating the development of domain generalization (DG) approaches that leverage source domain data to train models capable of generalizing to unseen target domains. In this paper, we are motivated by two major gaps in existing work on FL and DG: (1) the lack of formal mathematical analysis of DG objectives; and (2) DG research in FL being limited to the star-topology architecture. We develop Decentralized Federated Domain Generalization with Style Sharing (StyleDDG), a decentralized DG algorithm which allows devices in a peer-to-peer network to achieve DG based on sharing style information inferred from their datasets. Additionally, we provide the first systematic approach to analyzing style-based DG training in decentralized networks. We cast existing centralized DG algorithms within our framework, and employ their formalisms to model StyleDDG. We then obtain analytical conditions under which convergence of StyleDDG can be guaranteed. Through experiments on popular DG datasets, we demonstrate that StyleDDG can obtain significant improvements in accuracy across target domains with minimal communication overhead compared to baseline decentralized gradient methods.
Index Terms:
Decentralized Federated Learning, Domain Generalization
@@ I Introduction
Federated learning (FL) enables multiple edge devices to collaboratively train a machine learning (ML) model using data stored locally on each device [1]. Existing conventional “star-topology” FL methods rely on a central server to conventional sent from the clients [2]. However, in decentralized networks, a central server may not always be available to conduct model aggregations [3]. Motivated by this, recent research has investigated decentralized federated learning (DFL), in which edge devices iteratively communicate their model parameters with their one-hop neighbors after completing their local model training [4]. While synchronization among the participating edge devices in conventional FL is readily guaranteed through the aggregation at the end of each training round, DFL algorithms need to incorporate a cooperative consensus strategy to their learning process so that devices gradually learn the same model [5].
@@
When training ML models via FL/DFL, it is typically assumed that the test data is drawn from the same probability distribution as the training data [6]. However, this assumption does not always hold; in many real-world applications, a distribution shift can occur between the training set (source) and the test set (target) [7]. For example, consider a network of autonomous vehicles [8] aiming to improve their image classification models using DFL. As these devices are driven mostly during the morning and the afternoon, they mostly contain images in daylight (source domains), while their goal is to train a generalized model capable of classifying images taken at night (target domain) as well. It has been shown that ML algorithms can drastically fail to achieve generalization to such unseen target domains without making any special adjustments to their training processes [9]. Domain generalization (DG) aims to mitigate this issue, with the goal of preventing performance degradation under such distribution shifts [6].
@@
Due to the practical importance of enabling DG in FL settings, recent research has focused on federated domain generalization (FDG) [10], where DG algorithms are designed to operate in a distributed manner across edge devices. FDG introduces further challenges compared to centralized DG, primarily due to the limited number of source domains accessible to each device [11]. In FDG, in addition to label heterogeneity across devices, source heterogeneity must be also considered, further complicating the goal of achieving DG [12]. For instance, in the previously mentioned example, one device may contain images collected in the morning, while another holds images taken in the afternoon, yet the constructed global model is expected to make reliable predictions on night-time images.
@@
Research objectives. A large class of DG algorithms are style-based, with style statistics estimated from the source dataset and explored for generalization to the target. These works, whether for centralized training [6] or star-topology FL [10], lack a concrete mathematical formulation of the objective function being optimized to train domain-invariant models. Furthermore, the studied DG algorithms in FL are constrained by their star-topology networks, and cannot be readily extended to arbitrary decentralized networks. In this paper, we present one of the first style-based decentralized FDG algorithms, called Decentralized Federated Domain Generalization with Style Sharing (StyleDDG), in which edge devices communicate style information of their local datasets with their one-hop neighbors so that collectively they can train ML models that generalize well to unseen domains. Our StyleDDG algorithm, illustrated in Fig. 1, is inspired by the work in [13], where the authors present an FDG algorithm called StableFDG. Different than [13], StyleDDG is a fully decentralized DG algorithm, and does not add any further oversampling or attention layers as in StableFDG. The computation overhead of StableFDG is at least doubled compared to its analogous models, due to its oversampling step and the added attention layer, and this in turn obstructs our understanding of its style exploration strategy in achieving DG performance. Lastly, our other main contribution is to provide the first formal modeling of style-based DG training, both for well-known DG methods and for StyleDDG.
@@
  • We present StyleDDG, one of the first fully decentralized domain generalization (DDG) algorithms. In our setup, devices in a decentralized system aim to achieve DG without having any information about the whole network graph, and can only communicate with their one-hop neighbors, unlike in star-topology FL. Specifically, each device in DDG trains an ML model locally using its local dataset, and then it exchanges (i) its ML model parameters and (ii) style statistics of its current batch with its neighbors (see Fig. 1-(a)). Afterward, each device employs the received style statistics from its neighbors to explore the style space more effectively for DG (see Fig. 1-(b)).
  • Among all style-based DG methods, for the first time in the literature, we provide a systematic approach to mathematically formulate the training objective function and optimization procedure. As the operations of style-based methods have commonalities, we first derive a set of concrete formulations for well-known style-based DG methods. These generalized formulations lay the groundwork for formal modeling of future methods in this emerging field, applicable to both centralized and distributed/federated network settings.
  • Exploiting the above-mentioned analytical framework, we present a formal modeling of the objective function of StyleDDG and conduct convergence analysis, enabling us to provide one of the first convergence guarantees among all style-based DG papers. Our results illustrate that under certain conditions on the employed ML model, the smoothness of local loss functions can be guaranteed for each device. Consequently, this ensures that all devices collectively converge to the same optimal model.
@@ II Related Work
Decentralized federated learning (DFL). A recent line of work has considered FL over non-star topologies. This includes semi-decentralized FL [14, 15], which groups the devices into several subnetworks, with devices conducting intra-subnet communications before exchanging their model parameters with the centralized server. Fog learning generalizes this to multi-layer hierarchies between the edge devices and the backbone servers [16]. DFL is at the extreme end of this spectrum, where communication is entirely peer-to-peer, with the goal of iteratively updating and exchanging models over the device graph to arrive at an optimal ML model [17, 18, 19, 20, 21]. Nevertheless, currently there does not exist any DFL algorithm which is tailored for DG. In this paper, we propose a DG methodology tailored to the DFL network setting, leading to one of the first style-based DDG algorithms in the literature.
@@
Domain generalization (DG). DG aims to achieve generalization to out-of-distribution data from unseen target domains using distinct source domain data for training (see [6, 9] for recent surveys). While various classes of ML methods can be considered DG – like meta-learning [22, 23], self-supervised learning [24], regularization [25] and domain alignment [26, 27] – our focus in this paper is on the family of feature augmentation, style-based methods for DG. In particular, algorithms like $\operatorname{MixStyle}$ [28] and $\operatorname{DSU}$ [29] achieve DG via feature-level augmentation, where the styles of images are explored beyond the distribution of styles that exist within the source domains. In the core of these algorithms, they employ $\operatorname{AdaIN}$ [30] (adaptive instance normalization) to explore new features/styles during training. Compared to other DG methods, an advantage of style-based methods is that they do not require the domain labels to be known [13].
@@
Despite the notable progress in this domain, a major drawback of current research on DG is the lack of concrete formulations for objective functions. Thus, convergence guarantees for these methods have remained elusive. We fill this gap by rigorously formulating the optimization procedures of prior well-known DG methods and our approach, and by providing a convergence analysis in decentralized network settings.
@@
Federated domain generalization (FDG). Data heterogeneity is a notable challenge in FL, as the distribution of data among devices are non-IID [31]. This challenge becomes even more prominent when the distribution of test data is different than the distribution of training data available in devices, motivating the development of FDG algorithms. In this direction, FedBN [32] proposed to not average the local batch normalization layers to alleviate the feature shift problem. In StableFDG [13], each device uses one of the styles of other devices received from the server, and extrapolates the region outside the existing style domains to achieve generalization.
@@
However, existing FDG methods [33, 34, 35, 36] are tailored for conventional star-topology FL settings. DFL introduces novel challenges due to the need to incorporate peer-to-peer consensus mechanisms. In this paper, we present one of the first decentralized domain generalization (DDG) algorithms. Furthermore, we are unaware of any style-based DG algorithm with concrete convergence analysis. Our work thus fills an important gap through our formal modeling of the DG objective functions, which not only helps us gain insights to the theoretical implications of our proposed methodology, but also lays the foundation for rigorous formulation and convergence analysis of future style-based DG methods [37, 38].
@@ III Style-Based Domain Generalization
We begin with preliminaries in Sec. III-A. Then in Sec. III-B we discuss the problem formulation. Finally, we provide a novel mathematical formulation for the objective functions of two centralized DG algorithms in Sec. III-C, which we later use in Sec. IV to formulate our StyleDDG algorithm.
@@ III-A Preliminaries
Figure 2: Illustration of style statistics. (a) A decentralized network with devices having access to data from only a single domain. (b) An overview of the ResNet18 model [39] as an example model at each device, where the number of channels and output size of each block are indicated. We visualize the blocks/layers that StyleDDG is applied to, which correspond to layers where we extract style statistics from. (c) Illustrates the outputs of a particular layer for all instances in the batch of size $B$. The style statistics that each device shares with its neighbors are obtained in the batch-level as shown. (d) For a given instance in the batch, the style statistics are calculated for each channel separately. (e) The $\mu$ and $\sigma$ of style statistics are computed based on the mean and standard deviation of the layer outputs.
@@ III-A1 Notation
We denote the model parameters of a neural network with dimension $n\in\mathbb{Z}^{+}$ by $\theta\in\mathbb{R}^{n}$. Let us partition the model $\theta$ to $L\in\mathbb{Z}^{+}$ partitions. We use the notation $\theta_{\ell}\subset\theta$ with $\ell=1,...,L$ as the sub-model with the model parameters of only partition $\ell$ of the model $\theta$. Thus, we will have $\theta=\begin{bmatrix}\theta_{1}^{T}&...&\theta_{L}^{T}\end{bmatrix}^{T}$. We focus much of our presentation on convolutional neural networks (CNNs), given their widespread use in DG methods over vision/image applications. As an example, Fig. 2-(b) shows how a ResNet model [39] can be partitioned by its convolutional blocks.
@@ III-A2 $\operatorname{AdaIN}$
Adaptive instance normalization ($\operatorname{AdaIN}$) [30] plays a key role in style-based DG methods by enabling style transfers. This technique removes the style information of a data sample and replaces it with the style of another sample. Specifically, the $\operatorname{AdaIN}$ process can be written as follows:
@@
where the tensor $x\in\mathbb{R}^{B\times C\times H\times W}$ is taken to be a batch of size $B$ at a specific layer of the CNN with a channel dimension of $C$, width $W$ and height $H$, and $x_{b,c,h,w}$ denotes the value at the indexed dimensions. Furthermore, $\mu_{t},\sigma_{t}\in\mathbb{R}^{B\times C}$ are the statistics of a target style to which we want to shift the style of $x$ to, and the instance-level mean and standard deviation $\mu(x),\sigma(x)\in\mathbb{R}^{B\times C}$ for each channel are calculated as
@@
where $\mu(x)=[\mu(x)_{b,c}]_{\begin{subarray}{c}1\leq b\leq B\\ 1\leq c\leq C\end{subarray}}$, $\sigma(x)=[\sigma(x)_{b,c}]_{\begin{subarray}{c}1\leq b\leq B\\ 1\leq c\leq C\end{subarray}}$, and $\eta>0$ is a small value for numerical stability. We note that the $\operatorname{AdaIN}$ function in Eq. (1) first normalizes the style of $x$ by calculating $(x-\mu(x))/\sigma(x)$, and then scales the result by $\sigma_{t}$ and shifts it by $\mu_{t}$ to move it to a desired target style.
@@ III-B Problem Formulation
We consider a decentralized device-to-device (D2D) network represented by a graph $\mathcal{G}=(\mathcal{M},\mathcal{E})$, where $\mathcal{M}=\{1,...,m\}$ is the set of devices/nodes, and $\mathcal{E}$ indicates the set of communication edges between the devices. We assume that the graph is connected, and define the set of one-hop neighbors of a device $i\in\mathcal{M}$ as $\mathcal{N}_{i}$. In DFL [12, 15], device $i$ maintains an ML model $\theta_{i}$, and engages in (i) local model updates and (ii) inter-device communications to train a globally optimal model. Also, each device $i$ contains a local dataset $\mathcal{D}_{i}=\{(x_{i},y_{i},d_{i})_{k}\}_{k=1}^{|\mathcal{D}_{i}|}$ where the $(x_{i},y_{i},d_{i})_{k}$ pair indicates the $k$th data point with input features $x_{i}$, label $y_{i}$ and, unique to the DG setting, domain $d_{i}$. For example, in the well-known PACS dataset [40], each domain corresponds to a visual style, namely, art painting, cartoon, photo, or sketch.
In our decentralized DG (DDG) setting, the goal is to train a domain-invariant global model $\theta$ over the device network $\mathcal{G}$ that generalizes well to unseen domains. In the PACS dataset example, a global model trained on art painting, cartoon, and photo domains should generalize well to the sketch domain, despite the absence of sketch data and a central model aggregation server during the training process.

### v4

@@ Abstract
Much of federated learning (FL) focuses on settings where the local dataset statistics remain the same between training and testing. However, this assumption often does not hold in practice due to distribution shifts, motivating the development of domain generalization (DG) approaches that leverage source domain data to train models capable of generalizing to unseen target domains. In this paper, we are motivated by two major gaps in existing work on FL and DG: (1) the lack of formal mathematical analysis of DG objectives; and (2) DG research in FL being limited to the star-topology architecture. We develop Decentralized Federated Domain Generalization with Style Sharing (StyleDDG), a decentralized DG algorithm that allows devices in a peer-to-peer network to achieve DG based on sharing style information inferred from their datasets. Additionally, we provide the first systematic approach to analyzing style-based DG training in decentralized networks. We cast existing centralized DG algorithms within our framework and employ their formalisms to model StyleDDG. We then obtain analytical conditions under which the convergence of StyleDDG can be guaranteed. Through experiments on popular DG datasets, we demonstrate that StyleDDG can obtain significant improvements in accuracy across target domains with minimal communication overhead compared to baselines.
@@ I Introduction
Federated learning (FL) enables multiple edge devices to collaboratively train a machine learning (ML) model using data stored locally on each device [kairouz2021advances]. Existing conventional “star-topology” FL methods rely on a central server to aggregate the models sent from the clients [mcmahan2017communication]. However, in decentralized networks, a central server may not always be available to perform model aggregations [brinton2025key]. Motivated by this, recent research has investigated decentralized federated learning (DFL), in which edge devices iteratively communicate their model parameters with their one-hop neighbors after completing their local model training [yuan2024decentralized]. Although synchronization among edge devices that participate in conventional FL is readily guaranteed as they overwrite their local models with the one receiver from the server at the end of each training round, DFL algorithms must incorporate a cooperative consensus strategy into their learning process so that devices gradually learn the same model [nedic2009distributed].
@@
When training ML models using FL/DFL, it is typically assumed that test data are drawn from the same probability distribution as training data [zhou2022domain]. However, this assumption does not always hold; in many real-world applications, a distribution shift can occur between the training set (source) and the test set (target) [ben2010theory]. For example, consider a network of autonomous vehicles [filos2020can] with the aim of improving their image classification models using DFL. As these devices are driven mainly during the morning and afternoon, they mostly contain images in daylight (source domains), while their goal is to train a generalized model capable of classifying images taken at night (target domain) as well. ML algorithms have been shown to drastically fail in achieving generalization to such unseen target domains without making special adjustments to their training processes [wang2022generalizing]. Domain generalization (DG) aims to mitigate this issue, with the goal of preventing performance degradation under such distribution shifts [zhou2022domain].
@@
Due to the practical importance of enabling DG in FL settings, recent research has focused on federated domain generalization (FDG) [baibenchmarking], where DG algorithms are designed to operate in a distributed manner across edge devices. FDG introduces additional challenges compared to centralized DG, mainly due to the limited number of source domains accessible to each device [li2023federated]. In FDG, in addition to label heterogeneity between devices, source heterogeneity must also be considered, further complicating the goal of achieving DG [wang2024multi]. For instance, in the previously mentioned example, one device may contain images collected in the morning, while another holds images taken in the afternoon, yet the constructed global model is expected to make reliable predictions on night-time images.
@@
Research objectives. A large class of DG algorithms are style-based, with style statistics estimated from the source dataset and explored for generalization to the target. These works, whether for centralized training [zhou2022domain] or for star-topology FL [baibenchmarking], lack a concrete mathematical formulation of the objective function to be optimized in training domain-invariant models. Furthermore, the studied DG algorithms in FL are constrained by their star-topology networks, and cannot be readily extended to arbitrary decentralized networks. In this paper, we present one of the first style-based decentralized FDG algorithms, called Decentralized Federated Domain Generalization with Style Sharing (StyleDDG), in which edge devices communicate style information of their local datasets with their one-hop neighbors so that collectively they can train ML models that generalize well to unseen domains. Our StyleDDG algorithm, illustrated in Fig. 1, is inspired by the work in [park2024stablefdg], where the authors present an FDG algorithm called StableFDG. Unlike [park2024stablefdg], StyleDDG is a fully decentralized DG algorithm and does not add any additional oversampling or attention layers as in StableFDG. The computational overhead of StableFDG is at least doubled compared to its analogous models, due to its oversampling step and the added attention layer, and this in turn obstructs our understanding of its style exploration strategy in achieving DG performance. Lastly, our other main contribution is to provide the first formal modeling of style-based DG training, both for well-known DG methods and for StyleDDG.
@@
  • We present StyleDDG, one of the first fully decentralized domain generalization (DDG) algorithms. In our setup, devices in a decentralized system aim to achieve DG without having any information about the whole network graph, and can only communicate with their one-hop neighbors, unlike in star-topology FL. Specifically, each device in DDG trains an ML model locally using its local dataset, and then it exchanges (i) its ML model parameters and (ii) style statistics of its current batch with its neighbors (see Fig. 1-(a)). Afterward, each device employs style statistics received from its neighbors to explore the style space more effectively for DG (see Fig. 1-(b)).
  • Among all style-based DG methods, for the first time in the literature, we provide a systematic approach to mathematically formulate the training objective function and its optimization procedure. As the operations of style-based methods have commonalities, we first derive a set of concrete formulations for well-known style-based DG methods. These generalized formulations lay the groundwork for formal modeling of future methods in this emerging field, applicable to both centralized and distributed/federated network settings.
  • Exploiting the analytical framework mentioned above, we present a formal modeling of the objective function of StyleDDG and carry out its convergence analysis that enables us to provide one of the first convergence guarantees among all style-based DG work. Our results illustrate that under certain conditions on the ML model being employed, the smoothness of local loss functions can be guaranteed for each device. Consequently, this ensures that all devices collectively converge to the same optimal model.
@@ II Related Work
Decentralized federated learning (DFL). A recent line of work has considered FL over non-star topologies. This includes semi-decentralized FL [parasnis2023connectivity, chen2024taming], which groups devices into several subnetworks, with devices conducting intra-subnet communications before exchanging their model parameters with the centralized server. Fog learning generalizes this to multi-layer hierarchies between edge devices and backbone servers [hosseinalipour2020federated]. DFL is at the extreme end of this spectrum, where communication is entirely peer-to-peer, with the goal of iteratively updating and exchanging models on the network graph to arrive at an optimal ML model [liu2022general, zhang2022net, zhang2023novel, huang2024overlay]. Nevertheless, currently there does not exist any DFL algorithm which is tailored for DG. In this paper, we propose a DG methodology tailored to the DFL network setting, leading to one of the first style-based DDG algorithms in the literature.
@@
Domain generalization (DG). DG aims to achieve generalization to out-of-distribution data from unseen target domains using distinct source domain data for training (see [zhou2022domain, wang2022generalizing] for recent surveys). Although various classes of ML methods can be considered DG – like meta-learning [li2018learning, balaji2018metareg], self-supervised learning [carlucci2019domain], regularization [huang2020self] and domain alignment [ghifary2015domain, li2018domain] – our focus in this paper is on the family of feature augmentation, style-based methods for DG. In particular, algorithms such as $\operatorname{MixStyle}$ [zhou2021domain] and $\operatorname{DSU}$ [li2022uncertainty] achieve DG using feature-level augmentation, where image styles are explored beyond the distribution of styles that exist within the source domains. In the core of these algorithms, they employ $\operatorname{AdaIN}$ [huang2017arbitrary] (adaptive instance normalization) to explore new features/styles during training. Compared to other DG methods, an advantage of style-based methods is that they do not require the domain labels to be known [park2024stablefdg].
@@
Despite the notable progress in this domain, a major drawback of current research on DG is the lack of concrete formulations for objective functions. Thus, convergence guarantees for these methods have remained elusive. We fill this gap by rigorously formulating the optimization procedures of previous well-known DG methods and our approach, and by providing a convergence analysis in decentralized network settings.
@@
Federated domain generalization (FDG). Data heterogeneity is a core challenge in FL, as the distribution of data among devices is non-IID [li2020federated]. This challenge becomes more prominent when the distribution of test data is different from the distribution of training data on devices, motivating the development of FDG algorithms. In this direction, FedBN [lifedbn] proposed not to average the local batch normalization layers to alleviate the problem of feature shift. In StableFDG [park2024stablefdg], each device receives the style of one other participating device from the server, and uses it to extrapolate the region outside the existing style domains to achieve generalization.
@@
However, existing FDG methods [liu2021feddg, wu2021collaborative, chen2023federated, zhang2023federated] are tailored for conventional star-topology FL. DFL introduces novel challenges due to the need to incorporate peer-to-peer consensus mechanisms. In this paper, we present one of the first decentralized domain generalization (DDG) algorithms. Furthermore, we are unaware of any style-based DG algorithm with concrete convergence analysis. Our work thus fills an important gap through formal modeling of DG objective functions, which not only yields insight into the theoretical implications of our proposed methodology, but also lays the foundation for rigorous formulation and convergence analysis of future style-based DG methods [kang2022style, park2023test].
@@ III Style-Based Domain Generalization
We begin with our preliminaries in Sec. III-A, and discuss the problem formulation in Sec. III-B. Finally, we provide a novel mathematical formulation for the objective functions of two centralized DG algorithms in Sec. III-C, which we later use in Sec. IV to formulate our StyleDDG algorithm.
@@ III-A Preliminaries
Figure 2: Illustration of style statistics. (a) A decentralized network with devices having access to data from only a single domain. (b) An overview of the ResNet18 model [he2016deep] as an example model at each device, where the number of channels and output size of each block are indicated. We visualize the blocks/layers that StyleDDG is applied to, which correspond to layers where we extract style statistics from. (c) Illustrates the outputs of a particular layer for all instances in the batch of size $B$. The style statistics that each device shares with its neighbors are obtained in the batch-level as shown. (d) For a given instance in the batch, the style statistics are calculated for each channel separately. (e) The $\mu$ and $\sigma$ of style statistics are computed based on the mean and standard deviation of the layer outputs.
@@ III-A1 Notation
We denote the model parameters of a neural network with dimension $n\in\mathbb{Z}^{+}$ by $\theta\in\mathbb{R}^{n}$. Let us partition the model $\theta$ into $L\in\mathbb{Z}^{+}$ parts. We use the notation $\theta_{\ell}\subset\theta$ with $\ell=1,...,L$ as sub-model with the model parameters of partition $\ell$ of the model $\theta$. Thus, we will have $\theta=[\theta_{1}^{T},...,\theta_{L}^{T}]^{T}$. We focus much of our presentation on convolutional neural networks (CNNs), given their widespread use in DG methods over vision/image applications. As an example, Fig. 2-(b) shows how a ResNet model [he2016deep] can be partitioned by its convolutional blocks.
@@ III-A2 $\operatorname{AdaIN}$
Adaptive instance normalization ($\operatorname{AdaIN}$) [huang2017arbitrary] plays a key role in style-based DG methods by enabling style transfers. This technique removes the style information from a data sample and replaces it with the style of another sample. Specifically, the $\operatorname{AdaIN}$ process can be written as follows:
@@
where the tensor $x\in\mathbb{R}^{B\times C\times H\times W}$ is taken to be a batch of size $B$ in a specific CNN layer with a channel dimension of $C$, width $W$ and height $H$, and $x_{b,c,h,w}$ denotes the value in the indexed dimensions. Furthermore, $\mu_{t},\sigma_{t}\in\mathbb{R}^{B\times C}$ are the statistics of a target style to which we want to shift the style of $x$ to, and the instance-level mean and standard deviation $\mu(x),\sigma(x)\in\mathbb{R}^{B\times C}$ for each channel are calculated as
@@
where $\mu(x)=[\mu(x)_{b,c}]_{\begin{subarray}{c}1\leq b\leq B\\ 1\leq c\leq C\end{subarray}}$, $\sigma(x)=[\sigma(x)_{b,c}]_{\begin{subarray}{c}1\leq b\leq B\\ 1\leq c\leq C\end{subarray}}$, and $\eta>0$ is a small value for numerical stability. We note that the $\operatorname{AdaIN}$ function in Eq. (1) first normalizes the style of $x$ by calculating $(x-\mu(x))/\sigma(x)$, and then scales the result by $\sigma_{t}$ and shifts it by $\mu_{t}$ to move it to the desired target style.
@@ III-B Problem Formulation
We consider a decentralized device-to-device (D2D) network represented by a graph $\mathcal{G}=(\mathcal{M},\mathcal{E})$, where $\mathcal{M}=\{1,...,m\}$ is the set of devices/nodes, and $\mathcal{E}$ indicates the set of communication edges between the devices. We assume that the graph is connected and define the set of one-hop neighbors of a device $i\in\mathcal{M}$ as $\mathcal{N}_{i}$. In DFL [wang2024multi, chen2024taming], device $i$ maintains an ML model $\theta_{i}$, and engages in (i) local model updates and (ii) inter-device communications to train a globally optimal model. Moreover, each device $i$ contains a local dataset $\mathcal{D}_{i}=\{(x_{i},y_{i},d_{i})_{k}\}_{k=1}^{|\mathcal{D}_{i}|}$ where the pair $(x_{i},y_{i},d_{i})_{k}$ indicates the $k$th data point with input features $x_{i}$, label $y_{i}$, and unique to the DG setting, domain $d_{i}$. For example, in the well-known PACS dataset [li2017deeper], each domain corresponds to a visual style, namely, art painting, cartoon, photo, or sketch.
In our DDG setting, the goal is to train a global domain-invariant model $\theta$ on the network $\mathcal{G}$ that generalizes well to unseen domains. In the PACS dataset example, a global model trained in the art painting, cartoon, and photo domains should generalize well to the sketch domain, despite the absence of sketch data and a centralized aggregator during training.